# Long-horizon Locomotion and Manipulation on a Quadrupedal Robot with Large Language Models

### v1

@@ Abstract
We present a large language model (LLM) based system to empower quadrupedal robots with problem-solving abilities for long-horizon tasks beyond short-term motions. Long-horizon tasks for quadrupeds are challenging since they require both a high-level understanding of the semantics of the problem for task planning and a broad range of locomotion and manipulation skills to interact with the environment. Our system builds a high-level reasoning layer with large language models, which generates hybrid discrete-continuous plans as robot code from task descriptions. It comprises multiple LLM agents: a semantic planner for sketching a plan, a parameter calculator for predicting arguments in the plan, and a code generator to convert the plan into executable robot code. At the low level, we adopt reinforcement learning to train a set of motion planning and control skills to unleash the flexibility of quadrupeds for rich environment interactions. Our system is tested on long-horizon tasks that are infeasible to complete with one single skill. Simulation and real-world experiments show that it successfully figures out multi-step strategies and demonstrates non-trivial behaviors, including building tools or notifying a human for help.
@@ I Introduction
Figure 1: Overview of the hierarchical system for long-horizon loco-manipulation task. The system is built up from a reasoning layer for task decomposition (yellow) and a controlling layer for skill execution (purple). Given the language description of the long-horizon task (top), a cascade of LLM agents perform high-level task planning and generate function calls of parameterized robot skills. The controlling layer instantiates the mid-level motion planning and low-level controlling skills with RL.
We propose a hierarchical system to address the challenges in long-horizon loco-manipulation problems for quadrupeds as illustrated in Fig. 1. At the high level, we harness the planning ability of pretrained large language models (LLMs) for abstract reasoning over the long horizon. Leveraging the wealth of encoded common-sense knowledge, the LLM-based reasoning layer can directly predict high-level strategies without demonstrations or searching heuristics. To better ground to robotic problems, we design the reasoning module as a cascade of multiple LLM agents. A semantic planner sketches a discrete plan, while a parameter calculator predicts continuous parameters in the plan. Subsequently, a code generator encapsulates the plan into function calls for robot skills. At the low level, we train a spectrum of motion planning and control policies with model-free reinforcement learning, spanning from quadrupedal locomotion to bipedal manipulation.
Our system is tested with long-horizon loco-manipulation tasks such as turning off the lights before exiting the office, and delivering a package into a room with a closed door. The LLM-based reasoning module figures out strategic solutions including building stairs with appropriate heights to touch an originally unreachable button. Combined with versatile low-level motion planning and control policies, the system can address long-horizon tasks that require minutes of execution of multiple motion skills, demonstrating a success rate of over 70% in simulation and successful deployment in the real world.
@@ II Related Work
Quadrupedal locomotion and manipulation: Developing versatile motions over quadrupedal robots to make them work like real animals has attracted much research interest. A large body of works focuses on robust and agile locomotion skills that can walk on challenging terrains [9, 3, 10] and perform extreme parkour [11, 12]. Some recent works develop manipulation skills for quadrupeds to better interact with the environment, such as learning manipulation via locomotion to push heavy objects [13], manipulating with one or two legs while balancing on other legs [14, 7, 6], or manipulating with a gripper mounted on the robot [15, 16]. There are also works that learn mid-level motion planning policies over these low-level motion controllers for more complex skills, such as soccer manipulation [17, 18]. We focus on empowering quadrupeds with long-horizon problem-solving abilities beyond motion planning and control skills for specific tasks. The problems considered in this work require an abstract task-planning ability that reasons over the semantics. The desired strategy, such as building stairs under the wall so that the robot can climb up and reach the light switch, requires common-sense knowledge about the physical constraints of the environment and the robot’s capabilities and is challenging to discover from scratch.
@@
Recently, pretrained large language models [25, 26, 27] have demonstrated impressive reasoning abilities to plan over robot skills for solving long-horizon tasks. A line of works uses LLMs to decompose long-range tasks into step-by-step plans in natural language [28, 29, 30], and grounds the language plan to robot execution with language-conditioned control policies. With the development of code-completion LLMs, researchers have also considered using Pythonic code as the interface between LLM reasoning and robot skills [31, 32] since it can be more directly grounded to available robot APIs and allows more fine-grained behaviors using numerical arguments in function calls. We similarly leverage LLMs to generate robot code for completing long-horizon tasks that involve physical constraints in this work and separate semantic planning, numerical argument computing, and code writing to different LLM agents.
The most relevant work to us is RoboTool [33] which also reasons over long-range physical puzzles with multiple LLMs. However, RoboTool only leverages the basic quadrupedal locomotion skills, which essentially treats the quadruped robot as a mobile platform and does not fully utilize its high dimensionality. Our system additionally supports a more diverse set of low-level skills such as pitching up and bipedal manipulation, therefore can be applied to more complex scenarios.
@@ III-A High-level reasoning for long-horizon tasks with LLMs
To address the challenging reasoning and grounding problem, we design a cascade of three LLM agents with different focuses as the high-level reasoning module as shown in Fig. 2. A semantic planner focuses on task decomposition with constraints from physical affordances and robot capabilities and sketches a plan of primitive skills for the long-horizon problem. A parameter calculator figures out the precise arguments for the primitives in the plan. A code generator converts the plan and the arguments into one single piece of executable code that can strategically combine the available low-level skills to handle the long-horizon problem. We use GPT-4-turbo-preview [34] for all the LLM agents  [^1] We also tried other LLMs such as GPT-3.5 [26] and Kimichat [35]. GPT-3.5 struggles to generate feasible abstract plans due to its deficiency in reasoning. Kimichat can generate reasonable plans given the environment descriptions and constraints but fails to calculate the accurate parameters even when prompted with detailed instructions. .
@@ III-A1 Semantic planner
Figure 3: The decomposed plan generated by the semantic planner LLM. The plan unifies multiple control branches depending on the environment specification. The steps in the plan are either feasibility checks that could branch out different conditions, or sequences of primitive skills.
The output of the planner is shown in Fig. 3. It formulates the plan as a workflow with branches and jumps. Each step is either a feasibility check or an action step. A check step redirects the control flow to different steps depending on the environment configuration or raises an error if the task is deemed unsolvable given all the constraints. An action step invokes one or more parameterized skills to be executed sequentially, but the corresponding parameters may be incorrect or undefined, e.g., the predicted target position for the climbing skill in step 2 does not consider the size of the robot. The parameters are handled by a calculator described next.
@@ III-B1 Learning control policies for short-term motions
Quadrupedal locomotion is formulated as a velocity-tracking motion, and is trained with a multi-stage curriculum of increasingly challenging terrains to efficiently learn a robust policy. In the first stage, a walking policy is trained on a flat surface following the same reward specification of [38]. It is then transferred to the second stage to learn locomotion over static stairs where the stair height ranges from 0 to 0.35m. After the policy is capable of climbing up the highest stairs, it is transferred and continues to be trained in the final stage of environment where the stairs are built from multiple movable boxes. The policy is trained to track a linear velocity along a fixed direction across the boxes while avoiding moving the boxes at the same time. This final stage encourages the robot to learn a gentle and robust climbing strategy without stepping on the edges of the terrain.
Bipedal locomotion and manipulation To facilitate interaction with objects, we develop bipedal locomotion and manipulation policies that control the quadrupedal robot to reach target positions with its front legs while standing on its hind legs. We first train a bipedal locomotion policy that can pitch up from a four-leg standing pose, then track a linear bipedal walking velocity using the method in [39]. Afterward, we train a bipedal manipulation policy to reach target positions using one front leg with a summation of three categories of reward terms. The first category is position-tracking rewards that encourage the robot to move its left front toe to the desired position as close as possible. To reduce jitter around the desired position, the robot is also given a bonus reward linear to the number of consecutive steps during which the distance to the desired position is within a threshold. The second category shapes the robot base and the other three legs into an upright standing pose. The third category is regularization terms that penalize abrupt actions, large joint velocities, and dangerous collisions with the robot. The policy is trained from states in bipedal standing poses with the robot base pose and the joint positions sampled from a hand-crafted range. To transit from bipedal to quadrupedal standing poses, we also train a sit-down policy following the reward in [39].
@@ III-B3 Chaining RL skills
Directly executing different skills sequentially often leads to suboptimal behavior since the terminate state distribution of one policy may not perfectly match the initial state distribution of the next policy. For instance, the quadrupedal climbing policy is trained from initial states where the edges of all the stairs are parallel. However, the preceding box-pushing policy often terminates with the box deviating from the ideal pose right beneath the existing stairs. Executing the climbing skill from the out-of-distribution states will lead to missteps and falls, thus resulting in lower success rates as more skills are executed. To mitigate the impact of the preceding strategy’s termination state on the subsequent strategy, we collect a set of termination states of the preceding policy of each policy, then fine-tune the policy from a mixture of the original initial state distribution and the collected states. Note that we only run fine-tuning for the pair of policies that will be executed sequentially according to the plan generated by the reasoning layer to reduce the computation burden.
@@ IV Experiments
IV-A Experiment setup
We experiment with two benchmark tasks that require strategic high-level reasoning. One task is turning off the light with the button high beyond the reach of the robot. The desired solution is to first build up stairs from boxes, then climb on top of the stairs and stand up to press the button. The other task is delivering a package into a room with the door closed. The robot should ring the doorbell to ask a human inside the room to open the door, and then push the package. We generate robot code as illustrated in Fig. 2 for each task then execute the code in simulation or on the real robot.
To facilitate the successful deployment of RL policies to a real Xiaomi Cyberdog2 [40] robot, we add domain randomization to the simulation including joint friction, joint damping, mass displacement, motor strength, and observation noise to the object positions during training. We attach multiple AprilTags [41] to the surface of the objects and use a forward-view RealSense D430 camera on the robot for pose estimation. Since the tags frequently go out of the view of the cameras during deployment, we also mimic the phenomenon in simulation by freezing the relevant observations when the object is out of the camera’s field osf view. We also randomly freeze the observations to mimic the situation when the detection fails due to the motion blur of captured images.
IV-B Ablation studies
Ablation on the high-level reasoning layer: We first verify the proposed design of the high-level reasoning layer composed of a cascade of LLM agents in the simulation. We compare our method LLM (S+P+C) that leverages a semantic planner, a parameter calculator and a code generator with the following variants:
@@ IV-B Ablation studies
The performances are measured using (i) the overall success rate for completing the whole task, and (ii) a normalized distance metric defined specific to the semantic of each task. The light-switching task is considered successful when the distance between the front-left end-effector and the button is smaller than 0.03m in the direction perpendicular to the wall and 0.06m in the other two axises parallel to the wall. The normalized distance is calculated as the shortest distance between the robot and the button within one episode divided by the initial distance. The success condition for the package-delivery task is when the bounding box of the package is in the room. We also report the shortest distance between the package and a fixed spot behind the door within each episode normalized by the initial distance for this task.
We use a temperature of 0.2 to generate 3 runs of code for each method. The generated codes are evaluated for 100 trajectories in simulation. In each trajectory, we randomize the height and density of the boxes, the initial position and density of the package, the height of the button and doorbell, and the dynamics of the robot joints. For the light-switching task, the height of the lower stair will affect the necessity of pushing the box to construct the intermediate platform.
As reported in Table I, removing the parameter calculator and the semantic planner leads to a significant performance drop. Note that LLM (S+C) and LLM (C) fail to achieve any success in the light-switching task since they mostly get stuck at the first executed skill with incorrect arguments. The arguments generated without the parameter calculator fail to consider the bounding box of objects and their geometric relationship. As the code snippets in Fig. 4 show, LLM (S+C) sets the target position of climbing to the center of a stair rather than the top surface, and it decides to push the movable box exactly beneath the lower stair, which is physically infeasible. LLM (C) fails to consider basic physical concepts like volumes. For example, it calls a walking skill with the target position at the center of a box, which will collide with the box and make the later steps more difficult to succeed. Similarly, in the package-delivery task, the variants without the parameter calculator struggle to obtain the correct arguments. Some sampled code plans would even attempt to push the box into the locked room without ringing the bell.
Figure 4: The code snippets generated by different variants of the LLM-based reasoning layer in the light-switching task. Without the parameter calculator, the generated arguments by LLM (S+C) and LLM (C) fail to ground to the physical world. Without the semantic planner, the generated code from LLM (C) contains more unnecessary primitive actions.
Comparison with hierarchical RL: We also validate the LLM-based reasoning module by comparing it with a hierarchical RL (HRL) baseline. It trains a high-level RL policy that operates over the same set of low-level skills used in our system to function as the reasoning module. The high-level policy jointly predicts the discrete low-level skill ID and its continuous parameters every 0.1 seconds. We implement a two-head MLP with shared base layers and separate output layers for discrete and continuous predictions respectively. For the light-switching task, the policy is trained with a reward that encourages the robot to get its front-left end-effector closer to the light button, and for the package delivery task, it is trained with a reward regarding the distance between the package and the door. We early-terminate an episode when the robot falls or moves too far away from its initial position. The high-level policy is optimized with PPO [37]. As shown in Table I, HRL cannot outperform LLM-based methods. In the light-switching task, HRL learns a sub-optimal policy that consistently tries to climb straight up the stairs instead of using the box as a step when necessary. This results in many falls when trying to climb up the stairs. As for the package-delivery task, HRL only learns to perturb the package closer to the door a little bit with quadrupedal locomotion skills, and cannot discover how to unlock the door. They both fail to learn the correct solutions since the vast search space makes it almost impossible to explore the exact combination of skills without prior knowledge.
@@
TABLE I: Ablation studies on the reasoning layer. The mean and standard deviation of the success rate and a distance metric in two tasks across 3 seeds are reported. The composition of multiple LLM agents is critical to the good performance of the system. LLM-based reasoning performs better than an RL-based high-level policy.
@@
Light switching Package delivery
@@
Ablation on fine-tuning RL skills for smooth chaining: We study whether chaining RL skills via fine-tuning helps alleviate compounding errors by comparing our method with a variant without fine-tuning the low-level skills from augmented initial states. The two sets of skills are orchestrated with the same high-level code plans and are tested in the light-switching task. As reported in Table II, our method with additional fine-tuning performs significantly better in terms of the success rate and the relative distance metric. Since the two methods use the same high-level plans, it indicates that the proposed fine-tuning stage is critical to the robustness of the system when transiting between different skills.

### v2

@@ Abstract
We present a large language model (LLM) based system to empower quadrupedal robots with problem-solving abilities for long-horizon tasks beyond short-term motions. Long-horizon tasks for quadrupeds are challenging since they require both a high-level understanding of the semantics of the problem for task planning and a broad range of locomotion and manipulation skills to interact with the environment. Our system builds a high-level reasoning layer with large language models, which generates hybrid discrete-continuous plans as robot code from task descriptions. It comprises multiple LLM agents: a semantic planner for sketching a plan, a parameter calculator for predicting arguments in the plan, and a code generator to convert the plan into executable robot code. At the low level, we adopt reinforcement learning to train a set of motion planning and control skills to unleash the flexibility of quadrupeds for rich environment interactions. Our system is tested on long-horizon tasks that are infeasible to complete with one single skill. Simulation and real-world experiments show that it successfully figures out multi-step strategies and demonstrates non-trivial behaviors, including building tools or notifying a human for help. Demos are available on our project page: https://sites.google.com/view/long-horizon-robot.
@@ I Introduction
Figure 1: Overview of the hierarchical system for long-horizon loco-manipulation task. The system is built up from a reasoning layer for task decomposition and a controlling layer for skill execution. Given the language description of the long-horizon task, a cascade of LLM agents perform high-level task planning and generate function calls of parameterized robot skills. The controlling layer instantiates the mid-level motion planning and low-level controlling skills with RL.
We propose a hierarchical system to address the challenges in long-horizon loco-manipulation problems for quadrupeds as illustrated in Fig. 1. At the high level, we harness the planning ability of pretrained large language models (LLMs) for abstract reasoning over the long horizon. Leveraging the wealth of encoded common-sense knowledge, the LLM-based reasoning layer can directly predict high-level strategies without demonstrations or searching heuristics. To better ground to robotic problems, we design the reasoning module as a cascade of multiple LLM agents. A semantic planner sketches a discrete plan, while a parameter calculator predicts continuous parameters. Subsequently, a code generator encapsulates the plan into function calls for robot skills. At the low level, we train a spectrum of motion planning and control policies with model-free reinforcement learning, spanning from quadrupedal locomotion to bipedal manipulation.
Our contributions are as follows: We introduce a new reasoning architecture with multiple LLM agents which focuses on long-sequence robotic problems. Our system chains together both locomotion and manipulation skills, and enables more complex tasks that require nuanced physical interaction. Combined with the high-level reasoning capabilities of LLM, it can address complex long-horizon tasks that could span over 10 minutes of execution. We extensively validate our design components in four novel and challenging long-horizon tasks for quadrupedal robots including manipulating a light switch at a height, package delivery, bridge building, and taking an elevator. Our method outperforms all the baselines, demonstrating a success rate of over 70% in simulation and successful deployment in the real world.
@@ II Related Work
Quadrupedal locomotion and manipulation: Developing versatile motions over quadrupedal robots to make them work like real animals has attracted much research interest. A large body of work focuses on robust and agile locomotion skills that can walk on challenging terrains [9, 3, 10] and perform extreme parkour [11, 12]. Some recent works develop manipulation skills for quadrupeds to better interact with the environment, such as learning manipulation via locomotion to push heavy objects [13], manipulating with one or two legs while balancing on other legs [14, 7, 6], or manipulating with a gripper mounted on the robot [15, 16]. There are also works that learn mid-level motion planning policies over these low-level motion controllers for more complex skills, such as soccer manipulation [17, 18]. We focus on empowering quadrupeds with long-horizon problem-solving abilities beyond motion planning and control skills for specific tasks. The problems considered in this work require an abstract task-planning ability that reasons over the semantics. The desired strategy, such as building stairs alongside the wall so that the robot can climb up and reach the light switch, requires common-sense knowledge about the physical constraints of the environment and the robot’s capabilities and is challenging to discover from scratch.
@@
Recently, pretrained large language models [25, 26, 27] have demonstrated impressive reasoning abilities to plan over robot skills for solving long-horizon tasks. A line of work uses LLMs to decompose long-range tasks into step-by-step plans in natural language [28, 29, 30], and grounds the language plan to robot execution with language-conditioned control policies. With the development of code-completion LLMs, researchers have also considered using Pythonic code as the interface between LLM reasoning and robot skills [31, 32] since it can be more directly grounded to available robot APIs and allows more fine-grained behaviors using numerical arguments in function calls. We similarly leverage LLMs to generate robot code for completing long-horizon tasks that involve physical constraints in this work and separate semantic planning, numerical argument computing, and code writing to different LLM agents.
The most relevant work to us is RoboTool [33] which also reasons over long-range physical puzzles with multiple LLMs. However, RoboTool only leverages the basic quadrupedal locomotion skills, which essentially treats the quadruped robot as a mobile platform and does not fully utilize its high dimensionality. Furthermore, RoboTool is designed to generate plans specific to certain scene configurations, while our method can generate branching conditions to handle various scene configurations with the same piece of code. As we will see in Sec. IV-B, our method shows more robust planning performance.
@@ III-A High-level reasoning for long-horizon tasks with LLMs
To address the challenging reasoning and grounding problem, we design a cascade of three LLM agents with different focuses as the high-level reasoning module as shown in Fig. 2. A semantic planner focuses on task decomposition with constraints from physical affordances and robot capabilities and sketches a plan of primitive skills for the long-horizon problem. A parameter calculator figures out the precise arguments for the primitives in the plan. A code generator converts the plan and the arguments into one single piece of executable code that can strategically combine the available low-level skills to handle the long-horizon problem. We use GPT-4-turbo-preview [34] for all the LLM agents.
@@ III-A1 Semantic planner
The planner formulates the plan as a workflow with branches and jumps. Each step is either a feasibility check or an action step. A check step redirects the control flow to different steps depending on the environment configuration or raises an error if the task is deemed unsolvable given all the constraints. An action step invokes one or more parameterized skills to be executed sequentially, but the corresponding parameters may be incorrect or undefined, e.g., the predicted target position for the climbing skill in step 2 does not consider the size of the robot. The parameters are handled by a calculator described next.
@@ III-B1 Learning control policies for short-term motions
Quadrupedal locomotion is formulated as a velocity-tracking motion, and is trained with a multi-stage curriculum of increasingly challenging terrains to efficiently learn a robust policy. In the first stage, a walking policy is trained on a flat surface following the same reward specification of [37]. It is then transferred to the second stage to learn locomotion over static stairs where the stair height ranges from 0 to 0.35m. After the policy is capable of climbing up the highest stairs, it is transferred and continues to be trained in the final stage where the stairs are built from multiple movable boxes. The policy is trained to climb across the boxes while avoiding moving the boxes at the same time. This final stage encourages the robot to learn a gentle and robust climbing strategy without stepping on the edges of the terrain.
Figure 3: Versatile behaviors driven by the RL skill repertoire.
Bipedal locomotion and manipulation To facilitate interaction with objects, we develop bipedal locomotion and manipulation policies that control the quadrupedal robot to reach target positions with its front legs while standing on its hind legs. We first train a bipedal locomotion policy that can pitch up from a four-leg standing pose, then track a linear bipedal walking velocity using the method in [38]. Afterward, we train a bipedal manipulation policy to reach target positions using one front leg with a summation of three categories of reward terms. The first category is position-tracking rewards that encourage the robot to move its left front toe to the desired position as close as possible. To reduce jitter around the desired position, the robot is also given a bonus reward linear to the number of consecutive steps during which the distance to the desired position is within a threshold. The second category shapes the robot base and the other three legs into an upright standing pose. The third category is regularization terms that penalize abrupt actions, large joint velocities, and dangerous collisions. The policy is trained from states in bipedal standing poses with the robot base pose and the joint positions sampled from a hand-crafted range. To transit from bipedal to quadrupedal standing poses, we also train a sit-down policy following [38].
@@ III-B3 Chaining RL skills
Directly executing different skills sequentially often leads to suboptimal behavior since the terminal state distribution of one policy may not perfectly match the initial state distribution of the next policy.
For instance, the preceding box-pushing policy often terminates with the box deviating from the ideal pose right beneath the existing stairs. Executing the climbing skill from the out-of-distribution states will lead to missteps and falls, thus resulting in lower success rates as more skills are executed. To mitigate the impact of the preceding strategy’s termination state on the subsequent strategy, we collect a set of termination states of the preceding policy of each policy, then fine-tune the policy from a mixture of the original initial state distribution and the collected states. Note that we only run fine-tuning for the pair of policies that will be executed sequentially according to the plan generated by the reasoning layer to reduce the computation burden.
@@ IV Experiments
IV-A Simulation experiment setting
Figure 4: Total number of occurrences for different error types in generated code, with ten samples per variant. Spatial error refers to errors in perceiving spatial relations leading to miscalculations, constraint violation refers to ignoring the given constraints, and logical error refers to faulty logical reasoning. The lower part shows the execution of the 4 tasks in simulation.
IV-A1 Baselines
We first verify the proposed design of the high-level reasoning layer composed of a cascade of LLM agents in the simulation. We compare our method LLM (S+P+C) that leverages a semantic planner, a parameter calculator and a code generator with the following variants:
@@ IV-A1 Baselines
  • RoboTool [33] which is an LLM-based planning framework focusing on tool-use tasks.
  • Hierarchical RL [18]: We train a high-level RL policy that jointly predicts the discrete low-level skill ID and its continuous parameters every 0.1s.
To ensure a fair comparison of high-level reasoning with RoboTool, we constructed LLM agents following RoboTool’s structure and prompts, provided RoboTool with the same low-level skills and supplemented their original rules with additional information and computation rules from our method. We provided specific numerical parameters according to their environment description setting, which are not required by our LLM reasoning layer. For hierarchical RL, We implement a two-head MLP with separate output layers for discrete and continuous predictions respectively. The policy is optimized using PPO [36] and trained with a dense reward based on the aforementioned distance metric.
IV-A2 Benchmark tasks
We experiment with four benchmark tasks that require strategic high-level reasoning. (a) Light switching requires turning off the light with the button high beyond the reach of the robot. The robot needs to determine whether to build a step from boxes based on the height of the stairs, then climb on top of the stairs and stand up to press the button. (b) Package delivery is delivering a package into a room with the door closed. The robot should ring the doorbell to ask a human inside the room to open the door, and then push the package. (c) Bridge building requires the robot to build a bridge to reach the target platform. (d) Elevator ride, the robot needs to learn how to press the button to call the elevator, and select the correct floor button inside to reach the designated floor. This task provides the current and target floors, explicitly requiring the robot to decide whether to press the up or down button. We generate robot code as illustrated in Fig. 2 for each task then execute the code in simulation or on the real robot.
IV-A3 Evaluation metrics
@@
The performances are measured using (i) the overall Success rate for completing the whole task(examples of the desired strategies shown in Fig. 4), and (ii) a Normalized Distance metric defined specific to the semantic of each task. The normalized distance of light switching and elevator ride is calculated as the shortest distance between the robot’s toe and the button. We also report the shortest distance between the package and a fixed spot behind the door for package delivery, as well as the robot’s distance from the destination for bridge building. All the distance metrics are normalized by the initial distance.
We use a temperature of 0.2 to generate 3 runs of code for LLM-based methods and evaluate each code for 100 trajectories in simulation. In each trajectory, we randomize the sizes, initial poses, and physical properties of objects, as well as the dynamics of the robot joints. Since RoboTool is designed to generate code for specific parameterized scenarios, we regenerated the code for each trajectory.
@@
TABLE I: Ablation studies on the reasoning layer. The mean and standard deviation of the success rate and a distance metric in four tasks across 3 seeds are reported. The composition of multiple LLM agents is critical to the good performance of the system. LLM-based reasoning performs better than an RL-based high-level policy.
@@
Light switching Package delivery Bridge building Elevator ride
@@
[…]
Figure 5: Execution of LLM-generated long-term strategies for two benchmark tasks in the real world.
IV-B Ablation study analysis
As reported in Table I, removing the parameter calculator and the semantic planner leads to a significant performance drop. Note that LLM (S+C) and LLM (C) fail to achieve any success in most tasks since they mostly get stuck at the first executed skill with incorrect arguments. LLM (S+C) struggles to correctly consider the bounding box of objects and their geometric relationship. As shown in Fig. 4, the number of spatial miscalculations increases dramatically after the calculator is removed. For a more specific example, it sets the target position of climbing to the center of a stair rather than the top surface, and does not consider the robot’s height. It also decides to push the movable box exactly beneath the lower stair, which is physically infeasible. Without a semantic planner for preliminary analysis and planning, LLM (C) is more likely to make incorrect inferences and may even violate constraints, leading to ineffective or dangerous actions. In the bridge building task, it generated an incorrect execution sequence by attempting to push the box after climbing the low platform.
According to Table I, RoboTool demonstrates subpar performance across most tasks, which we attribute to three main reasons. The code generated by RoboTool lacks branching structures, limiting its adaptability to complex tasks. For instance, elevator ride requires determining whether to press the up or down button based on the current and target floors. Faced with a variety of possible combinations, RoboTool regenerates code for each new scenario, leading to inefficiency and instability. In contrast, our generated code uses an elegant branching structure that covers all possible scenarios for this task.ts. RoboTool is more likely to make mistakes in spatial calculation since it requires numerical values in the scene object description while simultaneously using variables in the prompts. As shown in Fig. 4, this mixed usage often introduced instability, leading to numerous errors in spatial relationship calculations. Furthermore, RoboTool summarizes the previous LLM agent output before passing it to the next. We found that these summaries often omit details necessary for effective task execution. Relevant prompts and code can be found from our website.
Comparison with hierarchical RL: As shown in Table I, HRL cannot outperform LLM-based methods. For example, in the package delivery task, HRL only learns to move the package slightly closer to the door but fails to unlock it. They both fail to learn the correct solutions since the vast search space makes it almost impossible to explore the exact combination of skills without prior knowledge.